# Evaluating Entity Retrieval in Electronic Health Records: a Semantic Gap Perspective

### v1

@@ Abstract.
Entity retrieval plays a crucial role in the utilization of Electronic Health Records (EHRs) and is applied across a wide range of clinical practices. However, a comprehensive evaluation of this task is lacking due to the absence of a public benchmark. In this paper, we propose the development and release of a novel benchmark for evaluating entity retrieval in EHRs, with a particular focus on the semantic gap issue. Using discharge summaries from the MIMIC-III dataset, we incorporate ICD codes and prescription labels associated with the notes as queries, and annotate relevance judgments using GPT-4. In total, we use 1,000 patient notes, generate 1,246 queries, and provide over 77,000 relevance annotations. To offer the first assessment of the semantic gap, we introduce a novel classification system for relevance matches. Leveraging GPT-4, we categorize each relevant pair into one of five categories: string, synonym, abbreviation, hyponym, and implication. Using the proposed benchmark, we evaluate several retrieval methods, including BM25, query expansion, and state-of-the-art dense retrievers. Our findings show that BM25 provides a strong baseline but struggles with semantic matches. Query expansion significantly improves performance, though it slightly reduces string match capabilities. Dense retrievers outperform traditional methods, particularly for semantic matches, and general-domain dense retrievers often surpass those trained specifically in the biomedical domain.
@@ 1. INTRODUCTION
Electronic Health Records (EHRs) are invaluable resources due to the rich patient information they contain (Osmani et al., 2017; Zhang et al., 2019). In clinical practice, retrieval is generally the first step to access the information in EHRs: physicians need to locate certain information for making clinical decisions (Ye et al., 2021; Pampari et al., 2018), and researchers need to search for specific criteria to find patients of interest (Li et al., 2021; Bouzillé et al., 2018). Such a process can be time-consuming (Arndt et al., 2017; Roman et al., 2017), and medical practitioners rely heavily on automatic EHR retrieval systems (Hanauer et al., 2015; Jackson et al., 2017).
@@
Based on CliniQ, we comprehensively benchmark various retrievers’ performances on EHR retrieval task, including BM25 (Robertson and Zaragoza, 2009), state-of-the-art dense retrievers of various sizes covering both general and medical domain, and the most capable proprietary embedding model by OpenAI. An overview of our benchmark results is provided in Figure 1 (b) and (c). In our experiments, the two settings present occasionally different model rankings. BM25 establishes a quite strong baseline, and the performance is further enhanced in Single-Patient Retrieval and recall@100 in Multi-Patient Retrieval through query expansion. Dense retrievers show consistent improvement with the increasing parameter size, and general domain retrievers outperform medical domain ones. With relevance annotations dissected by matching types, we first shed light on the semantic matching abilities of various retrievers in the context of EHR retrieval. We quantitatively reveal that the advantages of dense retrievers are mainly contributed by semantic matches. However, dense retrievers struggle in drug retrieval, where most queries are single-word and annotated by string match.
We believe this benchmark will be a valuable resource for the community, and our comprehensive analysis may point out future research directions.
@@ 2.1.1. Query set
Recently, using medical entities as queries has gained more and more research interest due to their versatility and their alignment with physicians’ practices (Yuan et al., 2020; Ruppel et al., 2020; Yang et al., 2011). Ping and Jinfa (2021) utilized $8$ cardiovascular disease terms, such as ”hypertension” and ”palpitation”, as queries; Yang et al. (2021) incorporated $20$ stroke-related concepts including diseases, symptoms, medications, and procedures; Shi et al. (2022) extracted $26$ disease mentions from radiology reports as the query set. However, all existing works used a quite small number of queries, with limited diversity, often focusing on specific information types or even specific diseases. The absence of a large-scale and diverse query set may hinder a comprehensive assessment of the models’ performance.
@@ 2.1.2. EHR corpus
Almost all existing evaluations on EHR retrieval use proprietary corpus (Voorhees, 2013; Wang et al., 2019; Thai et al., 2024). To the best of our knowledge, Myers et al. (2024) are the only ones attempting to utilize publicly available EHR corpus, MIMIC-III, to establish an EHR retrieval benchmark. However, the annotated dataset is not released, and a total of only $50$ patients are incorporated in the evaluation, potentially undermining the robustness of the evaluation. Besides, previous benchmarks generally adopt either Single-Patient (Ye and Fabbri, 2018; Ye et al., 2021; Gupta et al., 2024) or Multiple-Patient setting (Voorhees, 2013; Wang et al., 2019; Thai et al., 2024). None has assessed the model performances under different retrieval settings.
@@ 2.1.3. Relevance judgment
Traditionally, relevance judgment can be obtained solely from human experts (Voorhees, 2013; Wang et al., 2019; Yu et al., 2022), which is prohibitive to scale. In terms of positive labels, TREC annotated a total of $5,895$ relevant pairs in two years, and Wang et al. (2019) included $5,815$ positive annotations in their benchmark. Recently, researchers begin to explore automatic annotation methods to overcome the limitations of dataset scale. Thai et al. (2024) integrates Hypercube (Shekhar et al., 2023), a deterministic reasoning engine based on ontology, in the annotation pipeline and manages to achieve a Cohen’s Kappa efficient of $1$ with medical experts. However, the annotation process still relies on manually reading every patient record and extracting clinical facts regarding the queries, which is not generalizable and scalable. Myers et al. (2024) automatically annotated three specific types of mentions (diagnosis, surgeries, and antibiotics) using regular expressions and UMLS knowledge graph. Despite being totally untethered from intensive human labor, the annotation quality is strictly constrained by exact match, which is known to have a low recall (Koopman et al., 2016). This limitation may particularly underestimate more advanced methods such as dense retrieval (Thakur et al., 2021). In addition, none of existing benchmarks has incorporated detailed assessments on semantic match.
@@ 2.2.2. Patient cohort selection
In patient cohort selection, physicians or researchers aims to select a group of patients satisfying certain criteria for clinical researches or risk identification Sivarajkumar et al. (2024). This task can go beyond the domain of text retrieval since some criteria are related to structured patient characteristics such age and gender. For example, Cohort Retrieval Enhanced by Analysis of Text from Electronic Health Records (CREATE) is a system that performs cohort selection on both structured and unstructured data using the OMOP Common Data Model and Elasticsearch (Liu et al., 2020). Besides, the retrieval process relies largely on fixed vocabulary and Named Entitiy Recognition (NER) tools such as cTAKES (Savova et al., 2010). Facilitated by the data provided by TREC Medical Record tracks, the development of this area has been pushed relatively farther. More advanced methods including Siamese network (Xiao et al., 2020) and Transformer-based language model (Soni and Roberts, 2020) has been applied.
@@ 2.3. Dense Retrieval
In biomedical domain, the development of generalizable dense retrievers has been limited by the lack of high-quality paired data (Jin et al., 2023a). Prior to the advent of LLMs, Jin et al. (2023a) introduced MedCPT, which was trained on PubMed user logs and thus specialized in biomedical retrieval. At the time, MedCPT significantly outperformed general domain encoders in a wide range of biomedical retrieval tasks. Recently, leveraging powerful LLMs like GPT-3.5 and GPT-4 for data synthesizing, Xu et al. (2024) introduced BMRetriever, a series of LLM-based embedding model ($410$M, $2$B, and $7$B), further improving the performances on these tasks. However, the effectiveness of these models on the task of EHR retrieval remains unclear.
@@ 3.1. Corpus Pre-processing
We randomly sample $1,000$ discharge summary notes from MIMIC-III as our initial corpus. Following previous works (Mullenbach et al., 2018; Yuan et al., 2022), we perform basic data cleaning by removing masks in MIMIC, removing abundant white spaces, and lower-casing all tokens. We adopt the naive paragraph segmentation (Lanz and Pecina, 2024), simply splitting the notes into fixed length chunks with overlap. Formally, given the cleaned notes $\mathcal{N}=\{N_{1},N_{2},\dots\}$, we split each discharge summary $N_{i}$ into chunks $\mathcal{C}_{i}=\{c_{i}^{1},c_{i}^{2},\dots\}$, where each chunk $c_{i}^{j}$ has a length of $100$ words and two adjacent chunks $c_{i}^{j}$ and $c_{i}^{j+1}$ has 10-word overlap. Our final corpus consists of the union of all the chunks from separate notes, formally $\mathcal{C}=\bigcup_{i}\mathcal{C}_{i}$.
@@ 3.2. Query Curation
For ICD disease codes, we map all fine-sorted codes to their three-digit ancestors since the classifying system can be so detailed that the disease term becomes far more complicated than typical user queries, or even incomprehensible given the note. For example, we would map the code ”011.92”, Pulmonary tuberculosis, unspecified, bacteriological or histological examination unknown (at present), to code ”011”, Pulmonary tuberculosis. Besides, the ICD code term can contain phrases that are dependent on the coding system, such as ”other”, ”unspecified”, and ”not elsewhere classified”. To keep the queries self-contained, we manually remove such ambiguity.
For ICD procedure codes, which are less detailed and complex than the disease codes, we only apply the ambiguity removal process. For prescription codes, we utilize the National Drug Code (NDC) provided in MIMIC. We further remove information on usage, dosage, concentration, and formulation from the drug names using GPT-4o. The prompt we use is shown in Figure 2.
@@ 3.3. Relevance Judgment
The relevance judgment is given at chunk level, and we first perform a global exact match search for every query to annotate all string matches. As for semantic matches, we only annotate $\mathcal{Q}_{i}$ for each note $N_{i}$ due to cost constraints. To be specific, for each chunk $c_{i}^{j}$ in $N_{i}$, we prompt GPT-4o to generate relevance judgment for each query $q_{i}^{k}\in\mathcal{Q}_{i}$ (excluding string matched ones), and further classify the matching type into synonym, abbreviation, hyponym, and implication match. Formally, GPT-4o take $c_{i}^{j}$ and $\mathcal{Q}_{i}$ as input, and output a list of judgments $\mathcal{L}_{ij}=\{l_{ij}^{1},l_{ij}^{2},\dots\}$. Each $l_{i}^{k}$ takes value in =$\mathcal{R}=\{\textit{irrelevant},\textit{synonym},\textit{abbreviation},\textit{hyponym},\textit{implication}\}$, corresponding to the matching type (or irrelevance) of $q_{i}^{k}$. The prompt we use is shown in Figure 3. To enhance the annotation quality, we use Chain-of-Thought (CoT) prompt (Wei et al., 2022) for relevance judgment. Combining the two steps above, for each pair of chunk $c_{i}^{j}$ and query $q_{i}^{k}$, we obtain a label $l_{ij}^{k}\in\mathcal{R}\cup\{\textit{string}\}$.
@@
To evaluate the alignment of the automatic and manual annotations, we randomly sample $10$ notes and invite two senior M.D. candidates to annotate the relevance judgments and matching types. A chunk record is considered relevant if it contains information that is semantically related to the query and is of interest from the physician’s perspective when searching for the query. Disagreements are discussed to reach a ground truth label.
@@ 4.1. Experiment Settings
We do not offer a dissected benchmark in the Multi-Patient Retrieval setting due to the potential impacts of false negatives: the ICD and prescription codes in MIMIC are far from complete (van Aken et al., 2021), and thus some relevant codes may be missed in our annotation pipeline. Failing to remove all relevant chunks of other matching types may introduce significant bias to the measurements. On the other hand, in Single-Patient Retrieval, all pairs of each query and each chunk are annotated, including the irrelevant ones, so the benchmark can faithfully reflect the performances on each matching type.
@@ 4.2.3. RRF
RRF is a simple yet effective algorithm to combine the results from multiple retrievers to yield better performances (Cormack et al., 2009). For each document, the RRF score for merging $n$ retrievers’ results is calculated as $\sum_{i=1}^{n}\frac{1}{k+r_{i}}$, where $r_{i}$ is the rank of the document given by the $i$th retriever, and $k$ is hyperparameter. In our experiment, we set $k=60$ by convention. We combine sparse retriever augmented by query expansion and two best-performing dense retrievers, NV-Embed-v2 and text-embedding-3-large. In our experiments, the benefit of incorporating more retrievers is marginal, if any.
@@ 5.1. Dataset Statistics
The basic statistics of CliniQ are shown in Table 1. We report the total number of queries, chunks, and relevance judgments for Multi-Patient Retrieval, and report the average number per patient record for Single-Patient Retrieval. Starting from 1,000 discharge summaries in MIMIC-III, we split them into 16,550 chunks of 100 words, with an average of 16.6 chunks per document (Q1: 11.0; Q3: 21.0). The longest record in our corpus consists of 55 chunks. We collect 1,246 queries in total, comprising 329 diseases, 365 clinical procedures, and 552 drugs. In Single-Patient Retrieval, each record is associated with 27.9 queries on average (Q1: 18; Q3: 37; Max: 84), which results in a total of nearly 28k searches in this setting (within each record, we perform one search for each query associated).
@@
Human evaluation on a sample of 10 notes, 141 chunks reveals that our automatic annotations achieve a high alignment with medical experts on annotating relevance judgment and classifying matching types. Specifically, in terms of relevance judgment, considering only binary labels (relevant v.s. irrelevant), automatic annotations show a Cohen’s Kappa coefficient of 0.985 with human experts. With match types taken into consideration, where we formulate the annotation as a multi-class classification, our annotations achieve an accuracy of 99.5%.
Table 2. Performances of various retrieval methods on CliniQ.
@@
Table 3. Performances of various retrieval methods on Single-Patient Retrieval, dissected by match types.
@@
Table 4. Performances of various retrieval methods on different types of queries.

### v2

# CliniQ: A Multi-faceted Benchmark for Electronic Health Record Retrieval with Semantic Match Assessment

@@ Abstract.
Electronic Health Record (EHR) retrieval plays a pivotal role in various clinical tasks, but its development has been severely impeded by the lack of publicly available benchmarks. In this paper, we introduce a novel public EHR retrieval benchmark, CliniQ, to address this gap. We consider two retrieval settings: Single-Patient Retrieval and Multi-Patient Retrieval, reflecting various real-world scenarios. Single-Patient Retrieval focuses on finding relevant parts within a patient note, while Multi-Patient Retrieval involves retrieving EHRs from multiple patients. We build our benchmark upon $1,000$ discharge summary notes along with the ICD codes and prescription labels from MIMIC-III, and collect $1,246$ unique queries with $77,206$ relevance judgments by further leveraging powerful LLMs as annotators. Additionally, we include a novel assessment of the semantic gap issue in EHR retrieval by categorizing matching types into string match and four types of semantic matches. On our proposed benchmark, we conduct a comprehensive evaluation of various retrieval methods, ranging from conventional exact match to popular dense retrievers. Our experiments find that BM25 sets a strong baseline and performs competitively to the dense retrievers, and general domain dense retrievers surprisingly outperform those designed for the medical domain. In-depth analyses on various matching types reveal the strengths and drawbacks of different methods, enlightening the potential for targeted improvement. We believe that our benchmark will stimulate the research communities to advance EHR retrieval systems.
@@ 1. INTRODUCTION
Electronic Health Records (EHRs) are invaluable resources due to the rich patient information they contain (Osmani et al., 2017; Zhang et al., 2019). In clinical practice, retrieval is generally the first step to access the information in EHRs: physicians need to locate certain information for making clinical decisions (Ye et al., 2021; Pampari et al., 2018), and researchers need to search for specific criteria to find patients of interest (Li et al., 2021; Bouzillé et al., 2018). Such a process can be time-consuming (Arndt et al., 2017; Ying et al., 2025), and medical practitioners rely heavily on automatic EHR retrieval systems (Hanauer et al., 2015; Jackson et al., 2017).
@@
Based on CliniQ, we comprehensively benchmark various retrievers’ performance on EHR retrieval task, including BM25 (Robertson and Zaragoza, 2009), state-of-the-art dense retrievers of various sizes covering both general and medical domain, and the most capable proprietary embedding model by OpenAI. An overview of our benchmark results is provided in Figure 1 (b) and (c). In our experiments, the two settings present occasionally different model rankings. BM25 establishes a quite strong baseline, and the performance is further enhanced in Single-Patient Retrieval and recall@100 in Multi-Patient Retrieval through query expansion. Dense retrievers show consistent improvement with the increasing parameter size, and general domain retrievers outperform medical domain ones. With relevance annotations dissected by matching types, we first shed light on the semantic matching abilities of various retrievers in the context of EHR retrieval. We quantitatively reveal that the advantages of dense retrievers are mainly contributed by semantic matches, and among different types of semantic matches, implication match poses the greatest challenge for retrieval systems. Moreover, dense retrievers struggle in drug retrieval, where most queries are single-word and annotated by string match.
We believe this benchmark will be a valuable resource for the community, and our comprehensive analysis may point out future research directions. The benchmark is publicly available through Github (https://github.com/zhao-zy15/CliniQ) and HuggingFace (https://huggingface.co/datasets/THUMedInfo/CliniQ) [^1] Due to the credential requirement of MIMIC, we cannot redistribute the corpus directly. Rather, we release the hadm_ids of patients involved in our benchmark and the script to reproduce the corpus..
@@ 2.1.1. Query set
Recently, using medical entities as queries has gained more and more research interest due to their versatility and their alignment with physicians’ practices (Yuan et al., 2020; Ruppel et al., 2020; Yang et al., 2011). Ping and Jinfa (2021) utilized $8$ cardiovascular disease terms, such as ”hypertension” and ”palpitation”, as queries; Yang et al. (2021) incorporated $20$ stroke-related concepts including diseases, symptoms, medications, and procedures; Shi et al. (2022) extracted $26$ disease mentions from radiology reports as the query set. However, all existing works used a quite small number of queries, with limited diversity, often focusing on specific information types or even specific diseases. The absence of a large-scale and diverse query set may hinder a comprehensive assessment of the retrieval performance.
@@ 2.1.2. EHR corpus
Almost all existing evaluations on EHR retrieval use proprietary corpus (Voorhees, 2013; Wang et al., 2019; Thai et al., 2024). To the best of our knowledge, Myers et al. (2024) are the only ones attempting to utilize publicly available EHR corpus, MIMIC-III, to establish an EHR retrieval benchmark. However, the annotated dataset is not released, and a total of only $50$ patients are incorporated in the evaluation, potentially undermining the robustness of the evaluation. Besides, previous benchmarks generally adopt either Single-Patient (Ye and Fabbri, 2018; Ye et al., 2021; Gupta et al., 2024) or Multiple-Patient setting (Voorhees, 2013; Wang et al., 2019; Thai et al., 2024). None has assessed the model performance under different retrieval settings.
@@ 2.1.3. Relevance judgment
Traditionally, relevance judgment can be obtained solely from human experts (Voorhees, 2013; Wang et al., 2019; Yu et al., 2022), which is prohibitive to scale. In terms of positive labels, TREC annotated a total of $5,895$ relevant pairs in two years, and Wang et al. (2019) included $5,815$ positive annotations in their benchmark. Recently, researchers begin to explore automatic annotation methods to overcome the limitations of dataset scale. Thai et al. (2024) integrates Hypercube (Shekhar et al., 2023), a deterministic reasoning engine based on ontology, in the annotation pipeline and manages to achieve a Cohen’s Kappa efficient of $1$ with medical experts. However, the annotation process still relies on manually reading every patient record and extracting clinical facts regarding the queries, which is not generalizable and scalable. Myers et al. (2024) automatically annotated three specific types of mentions (diagnosis, surgeries, and antibiotics) using regular expressions and the UMLS knowledge graph. Despite being totally untethered from intensive human labor, the annotation quality is strictly constrained by exact match, which is known to have a low recall (Koopman et al., 2016). This limitation may particularly underestimate more advanced methods such as dense retrieval (Thakur et al., 2021). In addition, none of existing benchmarks has incorporated detailed assessments on semantic match.
@@ 2.2.2. Patient cohort selection
In patient cohort selection, physicians or researchers aims to select a group of patients satisfying certain criteria for clinical researches or risk identification (Sivarajkumar et al., 2024). This task can go beyond the domain of text retrieval since some criteria are related to structured patient characteristics such age and gender. For example, Cohort Retrieval Enhanced by Analysis of Text from Electronic Health Records (CREATE) is a system that performs cohort selection on both structured and unstructured data using the OMOP Common Data Model and Elasticsearch (Liu et al., 2020). Still, the retrieval process relies largely on fixed vocabulary and Named Entitiy Recognition (NER) tools such as cTAKES (Savova et al., 2010). Facilitated by the data provided by TREC Medical Record tracks, the development of this area has been pushed relatively farther. More advanced methods including Siamese network (Xiao et al., 2020) and Transformer-based language model (Soni and Roberts, 2020) has been applied.
@@ 2.3. Dense Retrieval
In biomedical domain, the development of generalizable dense retrievers has been limited by the lack of high-quality paired data (Jin et al., 2023a). Prior to the advent of LLMs, Jin et al. (2023a) introduced MedCPT, which was trained on PubMed user logs and thus specialized in biomedical retrieval. At the time, MedCPT significantly outperformed general domain encoders in a wide range of biomedical retrieval tasks. Recently, leveraging powerful LLMs like GPT-3.5 and GPT-4 for data synthesizing, Xu et al. (2024) introduced BMRetriever, a series of LLM-based embedding model ($410$M, $2$B, and $7$B), further improving the performance on these tasks. However, the effectiveness of these models on the task of EHR retrieval remains unclear.
@@ 3.1. Corpus Pre-processing
We randomly sample $1,000$ discharge summary notes from MIMIC-III as our initial corpus. Following previous works (Mullenbach et al., 2018; Yuan et al., 2022), we perform basic data cleaning by removing masks in MIMIC, removing abundant white spaces, and lower-casing all tokens. We further apply chunking to the notes to fit the context window constraints of dense retrievers and optimize the retrieval performance. Here we adopt the naive paragraph segmentation by simply splitting the notes into fixed length chunks with overlap since more sophisticated chunking strategies did not reveal consistent and significant improvement in previous work (Lanz and Pecina, 2024). Formally, given the cleaned notes $\mathcal{N}=\{N_{1},N_{2},\dots\}$, we split each discharge summary $N_{i}$ into chunks $\mathcal{C}_{i}=\{c_{i}^{1},c_{i}^{2},\dots\}$, where each chunk $c_{i}^{j}$ has a length of $100$ words and two adjacent chunks $c_{i}^{j}$ and $c_{i}^{j+1}$ has 10-word overlap. Our final corpus consists of the union of all the chunks from separate notes, formally $\mathcal{C}=\bigcup_{i}\mathcal{C}_{i}$.
@@ 3.2. Query Curation
For ICD disease codes, we map all fine-sorted codes to their three-digit ancestors since the classifying system can be so detailed that the disease term becomes far more complicated than typical user queries, or even incomprehensible given the note. For example, we would map the code ”011.92”, ”Pulmonary tuberculosis, unspecified, bacteriological or histological examination unknown (at present)”, to code ”011”, ”Pulmonary tuberculosis”. Besides, the ICD code terms can contain phrases that are dependent on the coding system, such as ”other”, ”unspecified”, and ”not elsewhere classified”. To keep the queries self-contained, we manually remove such phrases. For ICD procedure codes, which are less detailed and complex than the disease codes, we only process the codes by removing these ambiguous terms. For prescription codes, we utilize the National Drug Code (NDC) provided in MIMIC and remove information on usage, dosage, concentration, and formulation from the drug names using GPT-4o. The prompt we use is shown in Figure 2.
@@ 3.3. Relevance Judgment
The relevance judgment is given at chunk level, and we first perform a global exact match search for every query to annotate all string matches. As for semantic matches, we only annotate queries assigned to the patient in the coding systems $\mathcal{Q}_{i}$, rather than all queries $\mathcal{Q}$, for each note $N_{i}$ due to cost constraints. To be specific, for each chunk $c_{i}^{j}$ in $N_{i}$, we prompt GPT-4o [^2] We strictly conform to the data usage agreement of MIMIC by using Azure OpenAI service with proper certification. to generate relevance judgment for each query $q_{i}^{k}\in\mathcal{Q}_{i}$ (excluding string matched ones), and further classify the matching type into synonym, abbreviation, hyponym, and implication match. The formal definitions of the match types can be found in the prompt shown in Figure 3. Formally, GPT-4o takes $c_{i}^{j}$ and $\mathcal{Q}_{i}$ as input, and output a list of judgments $\mathcal{L}_{ij}=\{l_{ij}^{1},l_{ij}^{2},\dots\}$. Each $l_{i}^{k}$ takes value in =$\mathcal{R}=\{\textit{irrelevant},\textit{synonym},\textit{abbreviation},\textit{hyponym},\textit{implication}\}$, corresponding to the matching type (or irrelevance) of $q_{i}^{k}$. To enhance the annotation quality, we use Chain-of-Thought (CoT) prompt (Wei et al., 2022) for relevance judgment. Combining the two steps above, for each patient note $N_{i}$, we obtain complete pairwise relevance annotations between queries $\mathcal{Q}_{i}$ and chunks $\mathcal{C}_{i}$, classified into five match types (string match and four types of semantic matches).
@@
To evaluate the quality of the automatic annotations, we randomly sample a subset of CliniQ and invite two senior M.D. candidates to annotate the relevance judgments and matching types. A chunk record is considered relevant if it contains information that is semantically related to the query and is of interest from the physician’s perspective when searching for the query. Disagreements are discussed to reach a ground truth label.
@@ 4.1. Experiment Settings
We do not offer a dissected benchmark in the Multi-Patient Retrieval setting due to the potential impacts of false negatives: the ICD and prescription codes in MIMIC are far from complete (van Aken et al., 2021), and thus some relevant codes may be missed in our annotation pipeline. Failing to remove all relevant chunks of other matching types may introduce significant bias to the measurements. On the other hand, in Single-Patient Retrieval, all pairs of each query and each chunk are annotated, including the irrelevant ones, so the benchmark can faithfully reflect the performance on each matching type.
@@ 4.2.3. RRF
RRF is a simple yet effective algorithm to combine the results from multiple retrievers to yield better performance (Cormack et al., 2009). For each document, the RRF score for merging $n$ retrievers’ results is calculated as $\sum_{i=1}^{n}\frac{1}{k+r_{i}}$, where $r_{i}$ is the rank of the document given by the $i$th retriever, and $k$ is hyperparameter. In our experiment, we set $k=60$ by convention. We combine sparse retriever augmented by query expansion and two best-performing dense retrievers, NV-Embed-v2 and text-embedding-3-large. In our experiments, the benefit of incorporating more retrievers is marginal, if any.
@@ 5.1. Dataset Statistics
The basic statistics of CliniQ are shown in Table 1. We report the total number of queries, chunks, and relevance judgments for Multi-Patient Retrieval, and report the average number per patient record for Single-Patient Retrieval. Starting from 1,000 discharge summaries in MIMIC-III, we split them into 16,550 chunks of 100 words, with an average of 16.6 chunks per document (Q1: 11.0; Q3: 21.0). The longest record in our corpus consists of 55 chunks. We collect 1,246 queries in total, comprising 329 diseases, 365 clinical procedures, and 552 drugs. The collected queries cover 50.2% of the three-digit ICD-9 disease codes (329/655) and 13.6% of the ICD-9 procedure codes (365/2679). The coverage of drug codes is inaccessible due to the lack of a standardized drug vocabulary [^3] The NDC system is not a well-controlled dictionary like ICD. It lacks a one-to-one mapping between drug concepts and unique identifiers.. In Single-Patient Retrieval, each record is associated with 27.9 queries on average (Q1: 18; Q3: 37; Max: 84), which results in a total of nearly 28k searches in this setting (within each record, we perform one search for each query associated).
@@
Human evaluation is conducted on a sample of 141 chunks with 5,221 automatic annotations, derived from 10 randomly chosen patient notes. The subset for human evaluation covers 181 unique queries (about 10% of all queries in CliniQ) and is thus representative of the overall dataset quality. Table 2 presents the inter-annotator agreement between GPT-4o’s outputs, two human annotators, and the ground truth. The evaluation demonstrates that our automatic annotations achieve strong concordance with the ground truth, performing comparably to medical experts in both relevance judgment and matching type classification. The primary errors observed in both automated and manual annotations stem from over-implication: specifically, the assignment of implication labels to tenuously relevant cases.
Table 2. Human Evaluation Results. For relevance annotations, we only consider binary labels (relevant v.s. irrelevant) and report the Cohen’s Kappa coefficient. For match types, we formulate the annotation as a multi-class classification and use accuracy metric.
GPT-4o Annotator 1 Annotator 2
Relevance 0.985 0.989 0.992
Match Types 0.995 0.998 0.996
Table 3. Performance of various retrieval methods on CliniQ.
@@
Table 4. Performance of various retrieval methods on Single-Patient Retrieval, dissected by match types.
@@
Table 5. Performance of various retrieval methods on different types of queries.